### v1

@@ Abstract
OpenStreetMap (OSM), an online and versatile source of volunteered geographic information (VGI), is widely used for human self-localization by matching nearby visual observations with vectorized map data. However, due to the divergence in modalities and views, image-to-OSM (I2O) matching and localization remain challenging for robots, preventing the full utilization of VGI data in the unmanned ground vehicles and logistic industry. Inspired by the fact that the human brain relies on geometric and semantic understanding of sensory information for spatial localization tasks (Hermer and Spelke, 1994; Epstein and Vass, 2014), we propose the OSMLoc in this paper. OSMLoc is a brain-inspired single-image visual localization method with semantic and geometric guidance to improve accuracy, robustness, and generalization ability. First, we equip the OSMLoc with the visual foundational model to extract powerful image features. Second, a geometry-guided depth distribution adapter is proposed to bridge the monocular depth estimation and camera-to-BEV transform. Thirdly, the semantic embeddings from the OSM data are utilized as auxiliary guidance for image-to-OSM feature matching. To validate the proposed OSMLoc, we collect a worldwide cross-area and cross-condition (CC) benchmark for extensive evaluation. Experiments on the MGL dataset, CC validation benchmark, and KITTI dataset have demonstrated the superiority of our method. Code, pre-trained models, CC validation benchmark, and additional results are available on: https://github.com/WHU-USI3DV/OSMLoc.
@@
A brain-inspired framework, OSMLoc, which jointly learns the geometry, semantics, and pose for image-to-OpenStreetMap (I2O) localization;
A geometry-guided camera-to-BEV transformation module with depth prior from the foundation model;
An auxiliary semantic alignment task to improve the scene understanding capability of the image model;
A cross-area and cross-condition validation benchmark for extensive evaluation;
@@ 1 Introduction
Map Construction and positioning are the basics of seamless navigation for logistic robots and unmanned ground vehicles (UGV). Most existing industrial solutions use the geo-referenced images (Arandjelovic et al., 2016; Cheng et al., 2018; Hausler et al., 2021; Zhu et al., 2023) or pre-built 3D point cloud maps  (Yu et al., 2021; Chen et al., 2022; Shi et al., 2022; Zou et al., 2023) as the reference map. However, constructing geo-referenced images or point cloud maps is expensive and difficult to update. Since the beginning of the Internet era, volunteering geo-informatics (VGI) has emerged as a new paradigm that creates, assembles, and disseminates geographic data by crowd-sourcing users instead of specific experts  (Wang et al., 2024). With over 10 million registered contributors (OSM_Foundation, 2024), the OpenStreetMap (OSM) provides detailed and up-to-date data about stationary objects throughout the world, including infrastructure and other aspects of the built environment, points of interest, land use and land cover classifications, and topography (Fan et al., 2014; Vargas-Munoz et al., 2020). With rich geographic and semantic information, humans can easily localize themselves by comparing the surrounding views with the OSM web map. However, the image-to-OSM (I2O) localization process remains challenging for robots due to differing modalities and perspectives, which limits the full potential of VGI data in the logistics industry and hinder the development of navigation with the crowd-sourcing map.
@@
Due to above mentioned difficulties, research on I2O visual localization is scarce. Early I2O localization approaches, such as Samano et al. (2020) localize along the pre-defined routes, while Zhou et al. (2021) integrates deep image features into the Monte Carlo framework to improve the representation ability of the observation model. OrienterNet (Sarlin et al., 2023) is the first end-to-end single-image I2O localization approach, which lifts the front-view image features to the birds-eye-view (BEV) and estimates the pose by dense matching. MaplocNet (Wu et al., 2024) proposed a coarse-to-fine registration pipeline for I2O localization and introduced the semantic labels from the autonomous driving datasets as auxiliary supervision. Although existing methods have achieved impressive results, we highlight that two key challenges remain unsolved: 1) effective modular transformation of visual observations for cross-modality I2O matching; 2) explainable localization and strong generalization across diverse environments worldwide.
@@
Figure 1: Image-to-OpenstreetMap localization aims to estimate the 3 degrees of freedom (3-DoF) camera pose. (a) shows the core idea of our method, which integrates geometry and semantic guidance into the framework. (b) shows the global evaluation results.
@@
Inspired by how the human brain uses geometric and semantic understanding from sensory information and long-term spatial knowledge for visual localization task (Hermer and Spelke, 1994; Vass and Epstein, 2013; Epstein and Vass, 2014), this paper introduces the OSMLoc, a single image-based visual localization framework with semantic and geometric guidance. The core idea of OSMLoc is illustrated in Fig. 1(a). We inherit the fundamental model with rich prior knowledge, specifically, the DINOv2 encoder (Oquab et al., 2023) of the Depth Anything model (Yang et al., 2024) into our framework to construct powerful image features. A compact depth distribution adapter is proposed to incorporate geometric information from Depth Anything into camera-to-BEV transformation. To our knowledge, this is the first attempt to integrate the foundation model into the I2O visual localization framework. The semantic consistency between visual and OSM data aids the model’s learning in completing cross-modular matching. Moreover, existing same-area validation sets exhibit similar styles, with consistent building and road patterns within neighborhood blocks of the training sets. This similarity arises because the data is typically collected by the same users, on similar devices, and within close timeframes. This uniformity may introduce bias into the validation results. Therefore, we collect a novel cross-area and cross-condition (CC) validation benchmark for evaluation in diverse environments, as shown in Fig. 1(b). The CC validation benchmark contains thousands of images from Detroit, Munich, Taipei, and Brisbane with vast scenarios. Experiments on the Mapillary geo-localization (MGL) dataset, CC validation benchmark, and the KITTI dataset reveal that our method outperforms existing methods, particularly in unseen areas. Overall, the main contributions include:
@@
    We collect the cross-area and cross-condition (CC) validation benchmark with thousands of frames in Detroit, Munich, Taipei, and Brisbane, providing an extensive testbed for the I2O visual localization models.
@@ 2.1 Visual localization with expert data
The visual localization with point cloud has been studied for decades. Given a query image and pre-built point cloud map, the image-to-point cloud (I2P) visual localization task aims to estimate the transform parameters from the camera coordinate system to the point cloud coordinate system.  Sattler et al. (2011) proposed a direct I2P matching framework by assigning point features to visual words and constructing the 2D-3D correspondences by linear search.  Huitl et al. (2012) proposed a novel image and point cloud dataset called TumIndoor to explore the I2P visual localization task in indoor environments. Due to the significant computational and storage burden of point cloud, Cheng et al. (2016) proposed a point cloud simplification framework based on the K-Cover theory to improve efficiency. Recent approaches (Li et al., 2020; Chen et al., 2022; Shi et al., 2022) focus on learning-based methods to improve accuracy and robustness. Some other approaches (Li and Lee, 2021; Kang et al., 2024) assume the coarse location of the image is accessible and aim to estimate the precise 6-DoF pose. Although point cloud contains detailed spatial information for localization, the extraction, processing, and storage are costly and require specialized sensors and software, limiting their applicability in consumer-grade scenarios.
With wide coverage and easy access, aerial image-based visual localization methods advance rapidly. Early image-to-aerial image (I2A) approaches (Lin et al., 2013; Tian et al., 2017) simplify the localization task by framing it as a retrieval problem. They divide the aerial image into non-overlapping patches and search for the most similar one in the feature space. VIGOR (Zhu et al., 2021) pointed out that one-to-one image retrieval is not suitable for practical scenarios. Consequently, the authors proposed a novel dataset and a coarse-to-fine pipeline that first retrieves the coarse pose and then regresses the fine pose. Shi and Li (2022) argued that directly regressing the camera pose is very difficult and proposed a multi-scale pose optimization method. The method projects the aerial image onto the ground image plane via geometric projection and searches for the optimal camera pose by minimizing the feature difference. Recently, SNAP (Sarlin et al., 2024) solves the ground image and aerial image visual localization, semantic segmentation, and mapping with a unified framework. As previous approaches rely on accurate pose labels for supervision, Li et al. (2024) proposed an unsupervised framework to fully leverage unlabeled data. Ye et al. (2024) proposed a coarse-to-fine I2A localization pipeline for oblique-view imagery captured by unmanned aerial vehicles. Although aerial images are more compact than 3D point cloud maps, they are still expensive to capture, generally not free, and heavy to store at high resolution. Moreover, aerial images of the same location vary due to different times, seasons, weather conditions, and lighting, making the long-term visual localization task more challenging.
@@ 2.2 Visual localization with VGI
Unlike expert data, VGI data is created, assembled, and shared by individuals. Visual localization with the geo-referenced ground images from VGI platforms, such as Mapillary (Warburg et al., 2020) or Google Maps (Google, 2024) as the primary databases, is a common research area. Traditionally, the image-to-ground images (I2I) visual localization approaches follow a two-step pipeline: 1) finding the coarse location first; 2) estimating the precise pose subsequently. In the first step,  Arandjelovic et al. (2016); Li et al. (2023); Zhu et al. (2023) formulate the task as the visual place recognition problem, which retrieves the top candidates first, and then re-ranks the top-k candidates to improve the performance further. In the second step, matching-based methods (Dusmanu et al., 2019; Sarlin et al., 2020) construct the pixel-to-pixel correspondences first and then estimate the pose by solving for the homography, whereas matching-free methods (Kendall et al., 2015; Liu et al., 2020; Tang et al., 2024) directly regress the 6-DoF transformation parameters with a pose regressor.
OpenStreetMap-based visual localization approaches are most related to our method. The pioneering approach OpenStreetSLAM (Floros et al., 2013) integrates visual odometry with map information to improve the robustness and accuracy of the vehicle’s pose estimation. It extracts the street graph from the map and uses it as the geo-reference for the observation model in the Monto Carlo localization (MCL) framework. Samano et al. (2020) embeds the images and map tiles into low-dimension feature vectors and searches for the most similar OSM tile in the pre-defined routes. Zhou et al. (2021) combines the deep image features with a sequential MCL framework. Yan et al. (2019); Cho et al. (2022); Lee and Ryu (2024) utilize the point cloud as visual observation instead of the image for localization with the OSM data. OrienterNet (Sarlin et al., 2023) is the first end-to-end single-image I2O visual localization framework that lifts the image feature into the BEV space and estimates the camera pose via feature matching. However, the monocular depth estimation in the view transformation is unreliable and hard to generalize to unseen environments. MaplocNet (Wu et al., 2024) proposed a multi-task registration pipeline for simultaneously pose regression and semantic segmentation. The authors introduced the semantic labels from high-definition (HD) maps as auxiliary supervision. These labels are expertly annotated data and are not freely available. Unlike previous approaches, we propose a brain-inspired I2O visual localization framework that leverages geometric and semantic guidance from foundation models and OSM data, both are freely accessible and easily generalizable.
@@ 3 Our Approach: OSMLoc
[…]
Figure 2: Illustration of the OSM data rasterization.
The workflow of our method is illustrated in Fig. 3. Our OSMLoc first embeds the image $\mathcal{I}$ and OSM map tile $\mathcal{M}$ into high-dimensional feature space (Sec. 3.1), and then the depth distribution adapter distills the geometric prior from foundation model into depth predictions, while the camera-to-BEV transformation module (Sec. 3.2) lifts the image features to birds-eye-view. Lastly, the neural localization (Sec. 3.3) module estimates pose by dense feature matching. The semantic alignment is employed as an auxiliary task to improve the scene understanding ability of the image model. Pose classification loss $\mathcal{L}_{pose}$, depth distribution loss $\mathcal{L}_{depth}$ and semantic alignment loss $\mathcal{L}_{sem}$ are leveraged to supervise the network jointly.
@@ 3.1 Feature extraction with foundation model
For the image model, we utilize the DINOv2 (Oquab et al., 2023) of Depth Anything (Yang et al., 2024) as the image encoder. DINOv2 (Oquab et al., 2023) trained the vision transformer (ViT) (Dosovitskiy, 2020) based model over 142 M images for general vision perception tasks. Depth Anything (Yang et al., 2024) fine-tuned the pre-trained DINOv2 model on over 1.5 M labeled images and 62 M unlabeled images for the depth estimation task. As the Depth Anything model shows strong zero-shot relative depth estimation and semantic capability in various scenes, which is advantageous for the view transformation and neural localization, we adopt its DINOv2 model as our image encoder and freeze the pre-trained weights during training. Since Anyloc (Keetha et al., 2023) demonstrated that features from multiple stages of foundational models are beneficial for visual localization, we extract multi-stage features from the image encoder and feed them into a learnable feature pyramid network (FPN) decoder for progressive fusion. Finally, the output image feature $\boldsymbol{F}_{img}\in\mathbb{R}^{U\times V\times C}$ is bilinearly upsampled to the size $U\times V$ with $C$ channel.

### v3

@@ Abstract
OpenStreetMap (OSM), a rich and versatile source of volunteered geographic information (VGI), facilitates human self-localization and scene understanding by integrating nearby visual observations with vectorized map data. However, the disparity in modalities and perspectives poses a major challenge for effectively matching camera imagery with compact map representations, thereby limiting the full potential of VGI data in real-world localization applications. Inspired by the fact that the human brain relies on the fusion of geometric and semantic understanding for spatial localization tasks [1, 2], we propose the OSMLoc in this paper. OSMLoc is a brain-inspired visual localization approach based on first-person-view images against the OSM maps. It integrates semantic and geometric guidance to significantly improve accuracy, robustness, and generalization capability. First, we equip the OSMLoc with the visual foundational model to extract powerful image features. Second, a geometry-guided depth distribution adapter is proposed to bridge the monocular depth estimation and camera-to-BEV transform. Thirdly, the semantic embeddings from the OSM data are utilized as auxiliary guidance for image-to-OSM feature matching. To validate the proposed OSMLoc, we collect a worldwide cross-area and cross-condition (CC) benchmark for extensive evaluation. Experiments on the MGL dataset, CC validation benchmark, and KITTI dataset have demonstrated the superiority of our method. Code, pre-trained models, CC validation benchmark, and additional results are available at: https://github.com/WHU-USI3DV/OSMLoc.
@@
OSMLoc: a brain-inspired framework jointly fusing geometry and semantics for I2O localization.
Geometry-guided BEV transform with foundation model’s depth prior.
Auxiliary semantic alignment task boosts image model’s scene understanding.
Cross-area & cross-condition benchmark for extensive evaluation.
@@ 1 Introduction
The fusion of visual observation with geo-referenced maps forms the foundation for seamless navigation in logistic robots and unmanned ground vehicles (UGV) [3, 4, 5, 6]. Most existing industrial solutions use the geo-referenced images [7, 8, 9, 10] or pre-built 3D point cloud maps  [11, 12, 13, 14] as the reference map. However, constructing geo-referenced images or point cloud maps is expensive and difficult to update. Since the beginning of the Internet era, volunteering geo-informatics (VGI) has emerged as a new paradigm that creates, assembles, and disseminates geographic data by crowd-sourcing users instead of specific experts  [15]. Among VGI projects, OpenStreetMap (OSM) stands out as one of the most successful collaborative mapping initiatives, with over 10 million registered contributors [16]. OSM provides detailed and up-to-date data about stationary objects throughout the world, including infrastructure and other aspects of the built environment, points of interest, land use and land cover classifications, and topographical features [17, 18]. With its rich fusion of geographic and semantic content, humans can easily localize themselves by comparing the surrounding views with the OSM web maps. However, the image-to-OSM (I2O) localization process remains challenging for autonomous systems due to cross-modal discrepancies and drastic viewpoint variations, which limit the full potential of the OSM data in the logistics industry and hinder the development of navigation with the crowd-sourced map.
@@
Due to above mentioned difficulties, research on I2O visual localization is scarce. Early I2O localization approaches, such as Samano et al. [19] localize along the pre-defined routes, while Zhou et al. [20] integrates deep image features into the Monte Carlo framework to improve the representation ability of the observation model. OrienterNet [21] is the first end-to-end single-image I2O localization approach, which lifts the front-view image features to the birds-eye-view (BEV) and estimates the pose by dense matching. MaplocNet [22] proposed a coarse-to-fine registration pipeline for I2O localization and introduced the semantic labels from the autonomous driving datasets as auxiliary supervision. Although existing methods have achieved impressive results, we highlight that two key challenges remain unsolved: 1) effective modular transformation of visual observations for cross-modality fusion and I2O localization; 2) explainable localization and strong generalization across diverse environments worldwide.
@@
Figure 1: Image-to-OpenstreetMap localization aims to estimate the 3 degrees of freedom (3-DoF) camera pose. (a) shows the core idea of our method, which fuses geometric and semantic guidance into the framework. (b) shows the global evaluation results.
@@
Fortunately, the human navigation system provides profound inspiration for addressing these challenges. Hermer and Spelke [1] first demonstrated that even young children can utilize geometric cues from their surroundings for orientation and localization. Building on this, [2, 23] showed that the retrosplenial–hippocampal circuit in the brain supports robust localization through the integration of geometric and semantic information, a principle that directly motivates our framework. Further behavioral evidence from [24] indicates that humans construct cognitive maps by combining environmental geometry with landmark semantics. Building upon the foundation, we introduce the OSMLoc, a single image-based visual localization framework with fused geometric and semantic guidance. The core idea of OSMLoc is illustrated in Fig. 1(a). We inherit the fundamental model with rich prior knowledge, specifically, the DINOv2 encoder [25] of the Depth Anything model [26] into our framework to construct powerful image features. A compact depth distribution adapter is proposed to incorporate geometric information from Depth Anything into camera-to-BEV transformation. To our knowledge, this is the first attempt to integrate the foundation model into the I2O visual localization framework. The semantic consistency between visual and OSM data aids the model’s learning in completing cross-modular matching. Moreover, existing same-area validation sets exhibit similar styles, with consistent building and road patterns within neighborhood blocks of the training sets. This similarity arises from the fact the data is typically collected by the same users, on similar devices, and within close timeframes. Such uniformity may introduce bias into the validation results. Therefore, we collect a novel cross-area and cross-condition (CC) validation benchmark for evaluation in diverse environments, as shown in Fig. 1(b). The CC validation benchmark contains thousands of images from Detroit, Munich, Taipei, and Brisbane with vast scenarios. Experiments on the Mapillary geo-localization (MGL) dataset, CC validation benchmark, and the KITTI dataset reveal that our method outperforms existing methods, particularly in unseen areas. Overall, the main contributions include:
@@
    To evaluate generalization, we collect the cross-area and cross-condition (CC) validation benchmark with thousands of frames in Detroit, Munich, Taipei, and Brisbane, providing an extensive testbed for the I2O visual localization models.
@@ 2.1 Visual localization with expert data
The visual localization with point cloud has been studied for decades. Given a query image and pre-built point cloud map, the image-to-point cloud (I2P) visual localization task aims to estimate the transform parameters from the camera coordinate system to the point cloud coordinate system.  Sattler et al. [27] proposed a direct I2P matching framework by assigning point features to visual words and constructing the 2D-3D correspondences by linear search. Due to the significant computational and storage burden of point cloud, Cheng et al. [28] proposed a point cloud simplification framework based on the K-Cover theory to improve efficiency. Recent approaches [29, 13] focus on learning-based methods to improve accuracy and robustness. Although point cloud contains detailed spatial information for localization, the extraction, processing, and storage are costly and require specialized sensors and software, limiting their applicability in consumer-grade scenarios.
With broad coverage and easy accessibility, aerial image-based visual localization has seen rapid progress. Early image-to-aerial (I2A) methods [30, 31] treat localization as a retrieval task, dividing aerial images into patches and searching for the most similar one in feature space. VIGOR [4] noted the limitations of one-to-one retrieval in real-world scenarios and introduced a coarse-to-fine pipeline with a new dataset. SNAP [32] further unified ground-to-aerial localization, semantic segmentation, and mapping within a single framework. To reduce reliance on pose labels, Li et al. [33] proposed an unsupervised approach leveraging unlabeled data. While more compact than 3D maps, aerial images remain costly to obtain, often require paid access, and consume significant storage at high resolution. Additionally, variations across time, seasons, weather, and lighting conditions make long-term localization with aerial imagery particularly challenging.
@@ 2.2 Visual localization with VGI
Unlike expert data, VGI data is created, assembled, and shared by individuals. Visual localization with the geo-referenced ground images from VGI platforms like Mapillary [34] or KartaView [35] as the primary databases, is a common research area. Image-to-ground images (I2I) visual localization approaches follow a two-step pipeline: coarse location retrieval and precise pose estimation. For the first step,  [7, 36, 37, 38, 39] treat it as a place recognition task, which retrieves the top candidates and re-ranks the top-k candidates to refine results. For pose estimation, matching-based methods [40, 41] find pixel correspondences and solve for homography, whereas matching-free methods [42, 43, 44] directly regress 6-DoF poses with a neural pose regressor.
OSM-based visual localization approaches are most related to our method. The pioneering work OpenStreetSLAM [45] integrates visual odometry with OSM data to mitigate drift accumulated during vehicle motion. It extracts road line segments from the OSM tiles and uses them as geo-referenced constraints within a Monte Carlo Localization (MCL) framework. Samano et al. [19] re-formulated the I2O localization task as a global retrieval problem, embedding both query images and OSM map tiles into global descriptors for matching along predefined routes. Building on this formulation, Zhou et al. [20] incorporated these deep image/map features into an MCL framework to enable sequential localization. However, such methods are typically constrained to operate along predefined routes or limited areas, which restricts their applicability in large-scale, open-world scenarios. In addition, their localization accuracy remains relatively low due to the limited representation quality of learned features. More recently, Dong et al. [46] and Zhou et al. [47] combined camera images with LiDAR/Radar point clouds as visual observation to enhance the 3D geometric information. Other approaches [48, 49, 50, 51] utilize point clouds as the query for OSM-based localization. OrienterNet [21] is the first end-to-end, single-image I2O localization framework that lifts image features into the BEV space and estimates camera poses via feature matching. However, its monocular depth estimation for view transformation remains challenging, as the depth associated with each camera pixel is inherently ambiguous [52]. This limits both localization accuracy and generalization to unseen devices or scenes. MapLocNet [22] proposes a multi-task registration pipeline that performs simultaneous pose regression and semantic segmentation. Although it leverages drivable area labels from HD maps and building/POI labels from OSM as supervision, the HD maps are expert-curated and not widely available across diverse regions. Moreover, by ignoring diverse OSM semantic features, such as grass, paths, and road categories, the model overlooks rich geo-spatial context and may yield suboptimal performance. Since MapLocNet [22] relies on drivable area annotations from HD maps, which are expert-curated and not widely available across diverse regions, it requires fine-tuning on downstream datasets to achieve satisfactory performance. To address these limitations, we propose a novel I2O visual localization framework that integrates geometric and semantic cues extracted from foundation models and freely available OSM data. Our approach achieves competitive localization accuracy without relying on manual annotations or task-specific fine-tuning.
@@ 3 Our Approach: OSMLoc
Since our approach incorporates cross-view and cross-modality data, the image and OSM data require pre-processing before being fed into the framework. For the image $\mathcal{I}$, we rectify the roll and pitch angles to zero using the known gravity direction and ensure the principal axis is horizontal. The OSM data is stored in structured metadata and represents the areas, ways and nodes with the region of interest (ROI), as well as polylines and point of interest (POI). We categories the elements as in Fig. 2(a) and project the OSM data onto the x-y plane of the global coordinate system for rasterization. As shown in Fig. 2(b), we rasterize the areas, ways and nodes into a 3-channel grid map with a fixed sampling distance $\Delta s=50\,\text{cm/pixel}$, while preserving the semantic and geographic information. During training and evaluation, we crop map tile $\mathcal{M}$ of the rasterized OSM data centered around the location prior as input.
Figure 2: Illustration of the OSM pre-processing.
The workflow of our method is illustrated in Fig. 3. Our OSMLoc first embeds the image $\mathcal{I}$ and OSM map tile $\mathcal{M}$ into high-dimensional feature space (Sec 3.1), and then the depth distribution adapter distills the geometric prior from foundation model into depth predictions, while the camera-to-BEV transformation module (Sec 3.2) lifts the image features to birds-eye-view. Lastly, the neural localization (Sec 3.3) module estimates pose by dense feature matching. The semantic alignment is employed as an auxiliary task to improve the scene understanding ability of the image model. Pose loss $\mathcal{L}_{pose}$, disparity loss $\mathcal{L}_{dis}$ and semantic alignment loss $\mathcal{L}_{sem}$ are leveraged to supervise the network jointly.
@@ 3.1 Feature extraction with foundation model
For the image model, we adopt the DINOv2 [25] encoder from Depth Anything model (DAM) [26] as our backbone. DINOv2 [25] is a large-scale self-supervised vision transformer (ViT)[53] trained on 142 M diverse images, endowing the model with strong general-purpose visual representations that capture rich semantic concepts and structural cues. Building upon this, DAM fine-tuned DINOv2 on more than 1.5 M labeled and 62 M unlabeled images for monocular depth estimation, which further injects geometric priors into the encoder by explicitly learning relative depth relations across scenes.
Crucially, the resulting pre-trained encoder contains not only semantic information but also geometric depth cues, both of which are particularly important for localization. Semantic knowledge facilitates the recognition of place-related entities such as buildings, roads, and vegetation that remain stable across viewpoints, while geometric knowledge improves robustness to perspective changes and enables both view transformation and cross-modal alignment. To preserve these priors, we freeze the pre-trained encoder weights and extract multi-stage features that capture both low-level structures and high-level semantics. The final-layer features are fed into the DAM decoder to predict pseudo ground-truth normalized disparity $\hat{\boldsymbol{t}}$ for geometric-guided view transformation (described in Sec 3.2). As many works [54, 55] have shown that multi-level features from foundation models benefit localization, we progressively fuse the extracted multi-stage features using a feature pyramid network (FPN) decoder. The resulting image representation $\boldsymbol{F}_{img}\in\mathbb{R}^{U\times V\times C}$ is finally bilinearly upsampled to size $U\times V$ with $C$ channels.